OpenStudy (anonymous):

Let y(x) be the position of a particle at time x. Suppose that we know that the velocity of a particle satisfies the differential equation. y'(t) = ty with y(0)=1. We will try to give a reasonable method to approximate some of the positions of the particle. Integrate both sides to show that y(x)= ʃ ty(t)dt +1 from 0 to x

OpenStudy (dumbcow):

\[\frac{dy}{dt} = ty\] \[\frac{dy}{y} = t dt\] integrate \[\ln y = \frac{t^{2}}{2} + C\] \[\large y = k e^{t^{2}/2}\]